OpenStudy (anonymous):

How would I solve for -12w+39=-4w-17? Wouldn't subtract 39 from both side of the equation I don't know if I'm right and if I am what do I do after

OpenStudy (anonymous):

solve for W

OpenStudy (kendricklamar2014):

Step 1: Add 4w to both sides. \[−12w+39+4w=−4w−17+4w\] \[= −8w+39=−17\]

OpenStudy (kendricklamar2014):

Step 2: Subtract 39 from both sides. \[−8w+39−39=−17−39\] \[=\] \[−8w=−56\]

OpenStudy (kendricklamar2014):

Step 3: Divide both sides by -8. \[\frac{ -8w }{ -8 } = \frac{ -56 }{ -8 }\] \[ w = 7\]

OpenStudy (kendricklamar2014):

\[Answer: w = 7\]

OpenStudy (anonymous):

thank you

OpenStudy (kendricklamar2014):

You're Welcome

OpenStudy (anonymous):

It helped a lot

OpenStudy (kendricklamar2014):

Well im glad it did